OpenStudy (anonymous):

An emergency vehicle travels 10 miles at a speed of 50 miles per hour. How fast must the vehicle travel on the return trip if the round-trip travel time is to be 20 minutes?

OpenStudy (anonymous):

as the first step find out the time taken by the vechile to cover 10 miles

OpenStudy (anonymous):

how...?

OpenStudy (anonymous):

use the formula time = distance / speed

OpenStudy (anonymous):

time=10/50= 0.2= 0.2*60= 12 minutes... now what should be done @ Shizen

OpenStudy (anonymous):

round trip time covers the time to go and then come back so you need to subtract it from the roundtrip time which is given to be 20 convert that time into hour and then you need to use the formula speed = distance/time

OpenStudy (anonymous):

thanks... got my answer... bundle of thanks...

OpenStudy (anonymous):

your welcome :D